# LD-Scene: LLM-Guided Diffusion for Controllable Generation of Adversarial Safety-Critical Driving Scenarios

### v1

@@ Abstract
Ensuring the safety and robustness of autonomous driving systems necessitates a comprehensive evaluation in safety-critical scenarios. However, these safety-critical scenarios are rare and difficult to collect from real-world driving data, posing significant challenges to effectively assessing the performance of autonomous vehicles. Typical existing methods often suffer from limited controllability and lack user-friendliness, as extensive expert knowledge is essentially required. To address these challenges, we propose LD-Scene, a novel framework that integrates Large Language Models (LLMs) with Latent Diffusion Models (LDMs) for user-controllable adversarial scenario generation through natural language. Our approach comprises an LDM that captures realistic driving trajectory distributions and an LLM-based guidance module that translates user queries into adversarial loss functions, facilitating the generation of scenarios aligned with user queries. The guidance module integrates an LLM-based Chain-of-Thought (CoT) code generator and an LLM-based code debugger, enhancing the controllability and robustness in generating guidance functions. Extensive experiments conducted on the nuScenes dataset demonstrate that LD-Scene achieves state-of-the-art performance in generating realistic, diverse, and effective adversarial scenarios. Furthermore, our framework provides fine-grained control over adversarial behaviors, thereby facilitating more effective testing tailored to specific driving scenarios.
Index Terms:
Generative AI, diffusion models, Large Language Models, traffic simulation.
I Introduction
@@
With the continuous advancement of autonomous vehicle (AV) technologies, reliable testing is essential to verify the safety and robustness of self-driving systems [1, 2, 3]. However, real-world testing is not only expensive but also poses significant challenges in collecting safety-critical driving scenarios, which are crucial for assessing the performance of AVs. Specifically, safety-critical scenarios are characterized by high-risk interactions where AVs are exposed to unexpected maneuvers by other road users, which potentially lead to collisions [4, 5, 6]. As a more practical and efficient solution, the generation of safety-critical scenarios in simulation has become a widely adopted approach for driving performance evaluation.
@@
Existing approaches often employ simulators such as CARLA [7] and SUMO [8] to manually design safety-critical driving scenarios. However, these methods typically demand substantial domain expertise and often produce scenarios that lack realism. Recent studies have sought to address these limitations by leveraging large-scale driving datasets to learn realistic traffic models, subsequently generating safety-critical scenarios at test time through optimization-based techniques [9, 10, 11]. While these approaches improve realism to some extent, they remain limited in terms of generation efficiency and fine-grained behavioral control. More recently, diffusion models have demonstrated strong capabilities in modeling complex trajectory distributions through iterative denoising steps [12, 13]. Moreover, diffusion models have also been widely adopted in controllable generation tasks, such as text generation [14] and image synthesis [15], attracting considerable interest for their potential to generate realistic and controllable driving trajectories. Existing diffusion-based approaches employ various guidance mechanisms, such as reinforcement learning (RL)-based classifiers [16] and predefined objective functions [17, 18], to generate adversarial safety-critical scenarios. However, these approaches require retraining classifiers or redesigning objective functions with expert knowledge for different adversarial strategies, and these limit their flexibility and user-friendliness. With the above discussions, the key challenges in generating safety-critical scenarios lie in ensuring realism, achieving controllability over adversarial behaviors, and providing a user-friendly interface for scenario generation.
@@
Recent advancements in generative AI provide new solutions to deal with these challenges. In particular, Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding [19, 20] and code generation [21], making them particularly suitable for developing interfaces that allow users to specify adversarial scenarios in natural language. Meanwhile, Latent Diffusion Models (LDMs) [22] have gained attention for their ability to learn compact representations of complex driving scenarios, which facilitate more realistic and computationally efficient trajectory modeling. These advancements enhance both the expressiveness and efficiency of scenario generation. Inspired by these developments, integrating LLMs with LDMs presents a promising approach for generating realistic and controllable adversarial safety-critical driving scenarios while enabling intuitive and accessible user interaction.
@@
In this paper, we propose LD-Scene, a novel approach that integrates LLM-enhanced guidance with LDMs to enable user-controllable generation of adversarial safety-critical driving scenarios through natural language. As illustrated in Fig. 1, our LD-Scene framework consists of two main components: an LDM, which learns realistic driving trajectories, and an LLM-based guidance generation module, which translates user queries into a guidance loss function that perturbs the denoising process to generate adversarial driving scenarios. Specifically, both past information and future information are encoded into latent representations by a graph neural network (GNN)-based encoder, which captures vehicle interactions. The past latent serves as conditions for the denoising network to reconstruct a realistic future latent, ensuring the model learns plausible driving behaviors. The guidance generation module consists of a code generator and a code debugger. The Chain-of-Thought (CoT) reasoning process within the code generator enhances the controllability of adversarial scenario generation, allowing not only the specification of adversarial behaviors but also the adjustment of adversarial intensity. Meanwhile, the code debugger mitigates the instability inherent in LLM-based generation, thereby improving the robustness and effectiveness of the proposed approach.
@@
  • We propose LD-Scene, a novel framework that seamlessly integrates the LLM with an LDM to facilitate the generation of adversarial, safety-critical driving scenarios that are effortlessly controllable through natural language.
@@
  • Extensive experiments on the nuScenes dataset are performed, and the results attained demonstrate that LD-Scene outperforms baseline models in terms of adversariality, realism, and diversity, while also providing improved controllability over both the adversarial level and specific adversarial behaviors.
II Related Work
@@
II-A Safety-Critical Driving Scenario Generation
Simulating safety-critical scenarios is essential for comprehensive risk assessment of autonomous driving systems, especially since current AVs have been shown to perform well in typical driving conditions but remain under-tested in rare and hazardous cases [4]. Manual design of safety-critical scenarios, relying on expert knowledge and altering factors like actor locations and velocities, faces scalability issues and may yield implausible situations [23, 24, 25]. Recent studies delve into some specific parameterization spaces within original scenarios to pinpoint adversarial parameters using optimization-based methods [10, 11, 26, 27, 28, 29, 30]. Most works like AdvSim [10] directly perturb the standard trajectory space to produce adversarial trajectories, while adhering to constraints to maintain physical feasibility. Alternatively, some other works [27, 31, 9] perform parameter optimization within a condensed latent space. Strive [9] further incorporates GNNs [32] to generate traffic prior, while Mixsim [31] utilizes routes as priors both to enhance the plausibility of vehicle interactions. Nevertheless, these testing-time optimization methods encounter practical constraints that necessitate the iterative re-planning of a surrogate planner within the search loop, resulting in significant efficiency issues.
Other studies adopt adversarial policy models to address the sequential control of BVs in step-wise interactions. Building upon Proximal Policy Optimization (PPO) [33], NADE [34] trains background vehicles to execute adversarial maneuvers through simulation with a surrogate target model, offering flexibility and efficiency in scenario generation. However, this approach introduces complexity due to the curse of dimensionality and rarity. To alleviate this issue, D2RL [35] proposes a strategy to enhance information density by training neural networks with safety-critical data. They also selectively choose a critical state attacker to further reduce variance, sacrificing multi-vehicle trajectory rationality. The commendable efforts notwithstanding, adversarial policy training still necessitates a substantial number of interactions with the environment due to the inherent model complexity.
@@ II-A Safety-Critical Driving Scenario Generation
II-B Diffusion Models for Controllable Trajectory Generation
@@
Diffusion models [36, 22, 37] have emerged as a powerful class of generative models, attracting significant attention for their ability to generate realistic and controllable driving trajectories. Their capacity to effectively capture the complexity of real-world traffic patterns makes them well-suited for traffic simulation tasks [13]. Furthermore, guidance-based diffusion models improve controllability at test time while maintaining realism [38]. For example, Diffuser [39] and AdvDiffuser [16] leverage RL to train a reward function, which acts as a classifier to guide trajectory sampling. CTG [40] employs Signal Temporal Logic (STL) formulas as guidance for diffusion models to ensure compliance with specific rules, such as collision avoidance or goal-reaching. CTG++ [41] further incorporates language-based guidance, enabling more flexible rule enforcement. MotionDiffuser [42] proposes several differentiable cost functions as guidance, enabling the enforcement of both rules and physical constraints in the generated trajectories. In contrast, DiffScene [17] and Safe-Sim [18] focus on generating safety-critical traffic simulations by introducing safety-based guidance objectives. Furthermore, Safe-Sim [18] demonstrates the ability to generate controllable behaviors, such as specific collision types.
@@
Despite the successes achieved in the controllable generation of realistic driving trajectories, these prior works still face certain limitations in terms of user convenience and controllable generation efficiency. For example, RL-based guidance, such as AdvDiffuser [16], requires the training of different classifiers for various adversarial strategies. On the other hand, the controllability achieved through pre-designed objective-based guidance, such as Safe-Sim [18], often demands significant domain expertise, making it rather difficult for users without specialized knowledge in the field.
II-C Large Language Models for Traffic Simulation
@@
Recently, LLMs have represented a significant advancement in the field of autonomous driving [43, 44, 45], showcasing their potential in traffic simulation for generating realistic and controllable scenarios. This progress is largely driven by their capabilities in extensive knowledge storage, logical reasoning, and code generation [20, 21]. CTG++ [41] and [46] demonstrate LLM’s powerful code generation capabilities that can generate corresponding functions according to user requirements. Meanwhile, Scenediffuser [47] employs LLMs through few-shot prompting to convert scene constraints into structured Protobuf (Proto) representations, facilitating controllable scenario generation. LCTGen [48] employs LLMs as interpreters to transform textual queries into structured representations, which are subsequently combined with a transformer-based decoder to generate realistic traffic scenes. Building upon LCTGen, an enhanced method is introduced in [49], where the structured representations produced by LLMs incorporate interaction modeling, ultimately enhancing the generation of interactive traffic trajectories. On the other hand, ChatScene [50] utilizes an LLM-based agent to describe traffic scenes in natural language, subsequently converting these descriptions into executable Scenic code for scenario simulation within CARLA. Although these works have demonstrated a certain degree of controllability in traffic simulation, few studies have specifically focused on adversarial safety-critical scenario generation. In this context, it leaves an open and interesting problem to integrate LDMs with LLMs to achieve both realism and user-friendly controllability in the generation of safety-critical traffic scenarios.
III Methodology
@@ III-A Problem Formulation
Our work focuses on generating safety-critical driving scenarios for a given autonomous driving planner to facilitate a more effective and efficient evaluation of autonomous driving systems. Each scenario consists of $N$ agents, including an ego vehicle controlled by the planner $\pi$, while the future trajectories of the remaining $N-1$ vehicles are generated by our model. Among the $N-1$ vehicles, one is designated as the adversarial vehicle. The objective of our model is to encourage the adversarial vehicle to induce a collision with the ego vehicle while ensuring that the trajectories of the remaining non-adversarial vehicles remain realistic. Furthermore, our model incorporates LLMs to enable scenario customization based on user queries in natural language. This allows users to conveniently specify the adversarial vehicle, as well as control the collision type and adversarial collision severity.
@@
Similar to prior work [9, 16], a driving scenario $S$ consists of $N$ vehicle states and a map m that includes semantic layers for drivable areas and lanes. At any timestep $t$, the states of the $N$ agents are represented as $s_{t}=[s^{0}_{t},s^{1}_{t},...,s^{N-1}_{t}]$, where each agent state $s^{i}_{t}=(x^{i}_{t},y^{i}_{t},\theta^{i}_{t},v^{i}_{t})$ includes the 2D position, heading, and speed. The corresponding actions of the $N$ agents are represents as $a_{t}=[a^{0}_{t},a^{1}_{t},...,a^{N-1}_{t}]$, where each agent action $a^{i}_{t}=(\dot{v}_{i}^{t},\dot{\theta}_{i}^{t})$ representing the acceleration and yaw rate. The past trajectories, denoted as x, represent the historical states of all agents over the past $T_{hist}$ timesteps and are expressed as $\textit{{x}}=\{s_{t-T_{hist}},s_{t-T_{hist}+1},...,s_{t}\}$. The planner $\pi$ determines the future trajectory of the ego vehicle over a time horizon from $t$ to $t+T$. The planned future trajectory is denoted as $s^{0}_{t:t+T}=\pi(\textit{{m}},\textit{{x}})$, where $\pi$ processes historical state sequences and map feature to predict a sequence of future states.
@@
Our proposed LDM $g$, parameterized by $\theta$, is designed to generate multi-agent future trajectories in a driving scenario. The model consists of a pretrained encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given historical trajectory data and corresponding map information, the encoder $\mathcal{E}$ maps these inputs into a compact latent representation. A denoising diffusion process is then applied within this latent space, gradually refining the representation. Finally, the decoder $\mathcal{D}$ transforms the denoised latent representation into future trajectories of non-ego vehicles, denoted as $\bm{\tau}=\{s^{i}_{t:t+T}\}^{N-1}_{i=1}$. In the training stage, the model is trained on real-world driving data to learn realistic traffic behaviors. In the inference stage, we incorporate LLM-enhanced guidance to steer the denoising sampling process toward the generation of safety-critical scenarios, thereby ensuring both realism and controllability.
@@
III-B Latent Diffusion Models for Scenario Generation

### v2

@@ Abstract
Ensuring the safety and robustness of autonomous driving systems necessitates a comprehensive evaluation in safety-critical scenarios. However, these safety-critical scenarios are rare and difficult to collect from real-world driving data, posing significant challenges to effectively assessing the performance of autonomous vehicles. Typical existing methods often suffer from limited controllability and lack user-friendliness, as extensive expert knowledge is essentially required. To address these challenges, we propose LD-Scene, a novel framework that integrates Large Language Models (LLMs) with Latent Diffusion Models (LDMs) for user-controllable adversarial scenario generation through natural language. Our approach comprises an LDM that captures realistic driving trajectory distributions and an LLM-based guidance module that translates user queries into adversarial guidance functions, facilitating the generation of scenarios aligned with user queries. The guidance module integrates an LLM-based Chain-of-Thought (CoT) code generator and an LLM-based code debugger, enhancing the controllability and robustness in generating guidance functions. Extensive experiments conducted on the nuScenes dataset demonstrate that LD-Scene achieves state-of-the-art performance in generating realistic and effective adversarial scenarios. Furthermore, our framework provides fine-grained control over adversarial behaviors, thereby facilitating more effective testing tailored to specific driving scenarios.
keywords:
[…]
1 Introduction
@@
With the continuous advancement of autonomous vehicle (AV) technologies, reliable testing is essential to verify the safety and robustness of self-driving systems (Jiang et al., 2024c; Argui et al., 2024; Kang et al., 2019). However, real-world testing is not only expensive but also poses significant challenges in collecting safety-critical driving scenarios, which are crucial for assessing the performance of AVs. Specifically, safety-critical scenarios are characterized by high-risk interactions where AVs are exposed to unexpected maneuvers by other road users, which potentially lead to collisions (Ding et al., 2023; Wang et al., 2025; Zheng et al., 2025). As a more practical and efficient solution, the generation of safety-critical scenarios in simulation has become a widely adopted approach for driving performance evaluation.
@@
Existing approaches often employ simulators such as CARLA (Dosovitskiy et al., 2017) and SUMO (Lopez et al., 2018) to manually design safety-critical driving scenarios. However, these methods typically demand substantial domain expertise and often produce scenarios that lack realism. Recent studies have sought to address these limitations by leveraging large-scale driving datasets to learn realistic traffic models, subsequently generating safety-critical scenarios at test time through optimization-based techniques (Rempe et al., 2022; Wang et al., 2021; Zhang et al., 2022b). While these approaches improve realism to some extent, they remain limited in terms of generation efficiency and fine-grained behavioral control. More recently, diffusion models have demonstrated strong capabilities in modeling complex trajectory distributions through iterative denoising steps (Mao et al., 2023; Peng et al., 2024a). Moreover, diffusion models have also been widely adopted in controllable generation tasks, such as text generation (Li et al., 2022b) and image synthesis (Zheng et al., 2023), attracting considerable interest for their potential to generate realistic and controllable driving trajectories. Existing diffusion-based approaches employ various guidance mechanisms, such as reinforcement learning (RL)-based classifiers (Xie et al., 2024) and predefined objective functions (Xu et al., 2023; Chang et al., 2024), to generate adversarial safety-critical scenarios. However, these approaches require retraining classifiers or redesigning objective functions with expert knowledge for different adversarial strategies, and these limit their flexibility and user-friendliness. With the above discussions, the key challenges in generating safety-critical scenarios lie in ensuring realism, achieving controllability over adversarial behaviors, and providing a user-friendly interface for scenario generation.
@@
Recent advancements in generative AI provide new solutions to deal with these challenges. In particular, Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding (Achiam et al., 2023; Bubeck et al., 2023) and code generation (Guo et al., 2024), making them particularly suitable for developing interfaces that allow users to customize scenarios in natural language. For example, CTG++ (Zhong et al., 2023a) leverages language-based guidance to enable controllable generation of multi-agent traffic simulations. However, it lacks closed-loop evaluation of an ego planner and does not support reactive, adversarial safety-critical scenarios where an attacking agent challenges the planner’s decisions. In addition, the inherent instability of LLMs leads to frequent failures in generating valid guidance code, undermining the reliability of CTG++. Meanwhile, Latent Diffusion Models (LDMs) (Rombach et al., 2022) have gained attention for their ability to learn compact representations of complex driving scenarios, which facilitate more realistic and computationally efficient trajectory modeling. These advancements enhance both the expressiveness and efficiency of scenario generation. Inspired by these developments, integrating multi-agent LLM with LDMs presents a promising approach for generating realistic and controllable adversarial safety-critical driving scenarios while enabling intuitive and accessible user interaction.
@@
In this paper, we propose LD-Scene, a novel approach that integrates LLM-enhanced guidance with LDMs to enable user-controllable generation of adversarial safety-critical driving scenarios through natural language. As illustrated in Fig. 1, our LD-Scene framework consists of two main components: an LDM, which learns realistic driving trajectories, and an LLM-based guidance generation module, which translates user queries into a guidance loss function that perturbs the denoising process to generate adversarial driving scenarios. Specifically, both past information and future information are encoded into latent representations by a graph neural network (GNN)-based encoder, which captures vehicle interactions. The past latent serves as conditions for the denoising network to reconstruct a realistic future latent, ensuring the model learns plausible driving behaviors. The guidance generation module consists of a code generator and a code debugger. The Chain-of-Thought (CoT) reasoning process within the code generator enhances the controllability of adversarial scenario generation, allowing not only the specification of adversarial behaviors but also the adjustment of adversarial intensity. Meanwhile, the code debugger mitigates the instability inherent in LLM-based generation, thereby improving the robustness and effectiveness of the proposed approach. We conduct extensive closed-loop simulation experiments with an ego vehicle controlled by a rule-based planner to evaluate the effectiveness and realism of the generated scenarios. In addition, ablation studies validate the contribution of the debugger module to the stability of LD-Scene.
@@
  • We propose LD-Scene, a novel framework that seamlessly integrates the multi-agent LLM with an LDM to facilitate the generation of adversarial, safety-critical driving scenarios that are effortlessly controllable through natural language.
@@
  • Extensive experiments on the nuScenes dataset are conducted, and the results demonstrate that LD-Scene outperforms baseline models in terms of adversariality, realism, and stability, while also providing improved controllability over both the adversarial level and specific adversarial behaviors.
2 Related Work
@@
2.1 Safety-Critical Driving Scenario Generation
Simulating safety-critical scenarios is essential for comprehensive risk assessment of autonomous driving systems, especially since current AVs have been shown to perform well in typical driving conditions but remain under-tested in rare and hazardous cases (Ding et al., 2023). Manual design of safety-critical scenarios, relying on expert knowledge and altering factors like actor locations and velocities, faces scalability issues and may yield implausible situations (Scanlon et al., 2021; Li et al., 2022a; Zhang et al., 2022a). Recent studies delve into some specific parameterization spaces within original scenarios to pinpoint adversarial parameters using optimization-based methods (Wang et al., 2021; Zhang et al., 2022b; Abeysirigoonawardena et al., 2019; Ding et al., b; Hanselmann et al., 2022; Ding et al., a, 2021). Most works like AdvSim (Wang et al., 2021) directly perturb the standard trajectory space to produce adversarial trajectories, while adhering to constraints to maintain physical feasibility. Alternatively, some other works (Ding et al., b; Suo et al., 2023; Rempe et al., 2022) perform parameter optimization within a condensed latent space. Strive (Rempe et al., 2022) further incorporates GNNs (Scarselli et al., 2008) to generate traffic prior, while Mixsim (Suo et al., 2023) utilizes routes as priors both to enhance the plausibility of vehicle interactions. Nevertheless, these testing-time optimization methods encounter practical constraints that necessitate the iterative re-planning of a surrogate planner within the search loop, resulting in significant efficiency issues.
Other studies adopt adversarial policy models to address the sequential control of BVs in step-wise interactions. Building upon Proximal Policy Optimization (PPO) (Schulman et al., 2017), NADE (Feng et al., 2021) trains background vehicles to execute adversarial maneuvers through simulation with a surrogate target model, offering flexibility and efficiency in scenario generation. However, this approach introduces complexity due to the curse of dimensionality and rarity. To alleviate this issue, D2RL (Feng et al., 2023) proposes a strategy to enhance information density by training neural networks with safety-critical data. They also selectively choose a critical state attacker to further reduce variance, sacrificing multi-vehicle trajectory rationality. The commendable efforts notwithstanding, adversarial policy training still necessitates a substantial number of interactions with the environment due to the inherent model complexity.
@@ 2.1 Safety-Critical Driving Scenario Generation
2.2 Diffusion Models for Controllable Trajectory Generation
@@
Diffusion models (Ho et al., 2020; Rombach et al., 2022; Dhariwal and Nichol, 2021) have emerged as a powerful class of generative models, attracting significant attention for their ability to generate realistic and controllable driving trajectories. Their capacity to effectively capture the complexity of real-world traffic patterns makes them well-suited for traffic simulation tasks (Peng et al., 2024a). Furthermore, guidance-based diffusion models improve controllability at test time while maintaining realism (Jiang et al., 2024b). For example, Diffuser (Janner et al., 2022) and AdvDiffuser (Xie et al., 2024) leverage RL to train a reward function, which acts as a classifier to guide trajectory sampling. CTG (Zhong et al., 2023b) employs Signal Temporal Logic (STL) formulas as guidance for diffusion models to ensure compliance with specific rules, such as collision avoidance or goal-reaching. CTG++ (Zhong et al., 2023a) further incorporates language-based guidance, enabling more flexible rule enforcement. MotionDiffuser (Jiang et al., 2023) proposes several differentiable cost functions as guidance, enabling the enforcement of both rules and physical constraints in the generated trajectories. In contrast, DiffScene (Xu et al., 2023) and Safe-Sim (Chang et al., 2024) focus on generating safety-critical traffic simulations by introducing safety-based guidance objectives. Furthermore, Safe-Sim (Chang et al., 2024) demonstrates the ability to generate controllable behaviors, such as specific collision types.
@@
Despite the successes achieved in the controllable generation of realistic driving trajectories, these prior works still face certain limitations in terms of user convenience and controllable generation efficiency. For example, RL-based guidance, such as AdvDiffuser (Xie et al., 2024), requires the training of different classifiers for various adversarial strategies. On the other hand, the controllability achieved through pre-designed objective-based guidance, such as Safe-Sim (Chang et al., 2024), often demands significant domain expertise, making it rather difficult for users without specialized knowledge in the field.
2.3 Large Language Models for Traffic Simulation
@@
Recently, LLMs have represented a significant advancement in the field of autonomous driving (Yang et al., 2023; Peng et al., 2024b; Wen et al., 2023), showcasing their potential in traffic simulation for generating realistic and controllable scenarios. This progress is largely driven by their capabilities in extensive knowledge storage, logical reasoning, and code generation (Bubeck et al., 2023; Guo et al., 2024). CTG++ (Zhong et al., 2023a) and Guo et al. (2025) demonstrate LLM’s powerful code generation capabilities that can generate corresponding functions according to user requirements. Meanwhile, Scenediffuser (Jiang et al., 2024a) employs LLMs through few-shot prompting to convert scene constraints into structured Protobuf (Proto) representations, facilitating controllable scenario generation. LCTGen (Tan et al., 2023) employs LLMs as interpreters to transform textual queries into structured representations, which are subsequently combined with a transformer-based decoder to generate realistic traffic scenes. Building upon LCTGen, an enhanced method is introduced in Xia et al. (2024), where the structured representations produced by LLMs incorporate interaction modeling, ultimately enhancing the generation of interactive traffic trajectories. On the other hand, ChatScene (Zhang et al., 2024) utilizes an LLM-based agent to describe traffic scenes in natural language, subsequently converting these descriptions into executable Scenic code for scenario simulation within CARLA. Although these works have demonstrated a certain degree of controllability in traffic simulation, few studies have specifically focused on adversarial safety-critical scenario generation. In this context, it leaves an open and interesting problem to integrate LDMs with LLMs to achieve both realism and user-friendly controllability in the generation of safety-critical traffic scenarios.
3 Methodology
@@ 3.1 Problem Formulation
Our work focuses on generating safety-critical driving scenarios for a given autonomous driving planner to facilitate a more effective and efficient evaluation of autonomous driving systems. Each scenario consists of $N$ agents, including an ego vehicle controlled by the planner $\pi$, while the future trajectories of the remaining N1N-1 vehicles are generated by our model. Among the N1N-1 vehicles, one is designated as the adversarial vehicle. The objective of our model is to encourage the adversarial vehicle to induce a collision with the ego vehicle while ensuring that the trajectories of the remaining non-adversarial vehicles remain realistic. Furthermore, our model incorporates LLMs to enable scenario customization based on user queries in natural language. This allows users to conveniently specify the adversarial vehicle, as well as control the collision type and adversarial collision severity.
@@
Similar to prior work (Rempe et al., 2022; Xie et al., 2024), a driving scenario $S$ consists of $N$ vehicle states and a map m that includes semantic layers for drivable areas and lanes. At any timestep $t$, the states of the $N$ agents are represented as $s_{t}=[s^{0}_{t},s^{1}_{t},...,s^{N-1}_{t}]$, where each agent state $s^{i}_{t}=(x^{i}_{t},y^{i}_{t},\theta^{i}_{t},v^{i}_{t})$ includes the 2D position, heading, and speed. The corresponding actions of the $N$ agents are represents as $a_{t}=[a^{0}_{t},a^{1}_{t},...,a^{N-1}_{t}]$, where each agent action $a^{i}_{t}=(\dot{v}_{i}^{t},\dot{\theta}_{i}^{t})$ representing the acceleration and yaw rate. The past trajectories, denoted as x, represent the historical states of all agents over the past $T_{hist}$ timesteps and are expressed as $\textit{{x}}=\{s_{t-T_{hist}},s_{t-T_{hist}+1},...,s_{t}\}$. The planner $\pi$ determines the future trajectory of the ego vehicle over a time horizon from $t$ to $t+T$. The planned future trajectory is denoted as $s^{0}_{t:t+T}=\pi(\textit{{m}},\textit{{x}})$, where $\pi$ processes historical state sequences and map feature to predict a sequence of future states.
@@
Our proposed LDM $g$, parameterized by $\theta$, is designed to generate multi-agent future trajectories in a driving scenario. The model consists of a pretrained encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given historical trajectory data and corresponding map information, the encoder $\mathcal{E}$ maps these inputs into a compact latent representation. A denoising diffusion process is then applied within this latent space, gradually refining the representation. Finally, the decoder $\mathcal{D}$ transforms the denoised latent representation into future trajectories of non-ego vehicles, denoted as $\boldsymbol{\tau}=\{s^{i}_{t:t+T}\}^{N-1}_{i=1}$. In the training stage, the model is trained on real-world driving data to learn realistic traffic behaviors. In the inference stage, we incorporate LLM-enhanced guidance to steer the denoising sampling process toward the generation of safety-critical scenarios, thereby ensuring both realism and controllability.
@@
3.2 Latent Diffusion Models for Scenario Generation